OpenStudy (anonymous):

Show that a line through (0, 0) and (c, d) is perpendicular to a line through (0, 0) and (-d, c). Slope of first line = Slope of second line =

OpenStudy (ash2326):

johny find the slope of both the lines , let them be m1 and m2 if \[m1*m2=-1\] then the lines are perpendicular

OpenStudy (anonymous):

Two lines are perpendicular if the product of there slopes is -1

OpenStudy (anonymous):

i already knew that

OpenStudy (anonymous):

For the line through (0, 0) and (c, d) the slope is \( \frac {d}{c} \) and for the line through (0,0) and (-d.c) the slope is \(-\frac{c}{d} \) now if you take the product it will be -1 hence the two lines will be perpendicular.